OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\normalsize \text{calculate }\hspace{.33em}\\~\\ \cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

\[\cos(90-\theta) = \sin(\theta)\]

OpenStudy (perl):

oh that was a big hint :)

OpenStudy (perl):

may i expand on it ?

OpenStudy (perl):

First regroup the terms like so cos^(2) + cos(88)^2 + cos(4)^2 + cos(86)^2 + ...

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ =\sin^2 88^{\circ}+\sin^2 86^{\circ}+\sin^2 84^{\circ}+\cdot \cdot \cdot \cdot \cdot+\sin^2 2^{\circ}+\sin^2 0^{\circ}-------\color{red}{(1)} \hspace{.33em}\\~\\ \cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ =45-\sin^2 2^{\circ}+\sin^2 4^{\circ}+\sin^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\sin^2 90^{\circ}-------\color{red}{(2)} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (perl):

First regroup the terms like so cos^(2) + cos(88)^2 + cos(4)^2 + cos(86)^2 + ... = cos(2)^2 + sin(2)^2 + cos(4)^2 + sin(4)^2 + ... = 1 + 1 + ...

OpenStudy (perl):

cos^2 90 doesnt match with anything, but cos(90) is zero

OpenStudy (mathmath333):

in the method i did , what is wong

OpenStudy (mathmath333):

*wrong

OpenStudy (perl):

how did you get 45

OpenStudy (perl):

it looks like you went back and forth , changing all cosines to all sines, then back to all cosines

OpenStudy (mathmath333):

i applied the rule \(\large\tt \begin{align} \color{black}{1-\sin^2 \theta =\cos^2 \theta \hspace{.33em}\\~\\ }\end{align}\) and there are \(45\) terms overall

OpenStudy (mathmath333):

this it the corrected one \(\Huge \downarrow \) \(\large\tt \begin{align} \color{black}{\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ =\sin^2 88^{\circ}+\sin^2 86^{\circ}+\sin^2 84^{\circ}+\cdot \cdot \cdot \cdot \cdot+\sin^2 2^{\circ}+\sin^2 0^{\circ}-------\color{red}{(1)} \hspace{.33em}\\~\\ \cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ =45-(\sin^2 2^{\circ}+\sin^2 4^{\circ}+\sin^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\sin^2 90^{\circ})-------\color{red}{(2)} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{2(\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ}) \hspace{.33em}\\~\\ =\sin^2 88^{\circ}+\sin^2 86^{\circ}+\sin^2 84^{\circ}+\cdot \cdot \cdot \cdot +\sin^2 2^{\circ}+\sin^2 0^{\circ}+\hspace{.33em}\\~\\ 45-(\sin^2 2^{\circ}+\sin^2 4^{\circ}+\sin^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\sin^2 90^{\circ} \hspace{.33em}\\~\\ =45-\sin^2 90^{\circ} \hspace{.33em}\\~\\ 2(\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ})=44 \hspace{.33em}\\~\\ so~~\hspace{.33em}\\~\\ (\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ})=22 \hspace{.33em}\\~\\ }\end{align}\) is that right ?

OpenStudy (perl):

let me compare to my solution, one sec

OpenStudy (perl):

cos^2(2) + cos^2(4) + ... +cos^2 (88) + cos^2(90) =cos^2(2) + cos^2(4) + ... +cos^2 (88) + 0 =cos^2(2) + cos^2(4) + ... +cos^2 (88) =(cos^(2) + cos(88)^2) +(cos(4)^2 + cos(86)^2) + ... (cos^2(44)^2 + cos^2(46)) = (cos(2)^2 + sin(2)^2 ) + (cos(4)^2 + sin(4)^2) + ... (cos^2(44)^2 + sin^2(44)) = 1 + 1 + ... 1 = 22*1 = 22

OpenStudy (perl):

yes we got the same :)

OpenStudy (perl):

im a little confused about your lines , when you have ____ _ _ _ _ _ _

OpenStudy (perl):

you didnt square them

OpenStudy (mathmath333):

wait i forgot squaring lol

OpenStudy (perl):

you can put latex right into wolfram? thats cool

OpenStudy (perl):

yes i got the same , so you can use sigma notation

OpenStudy (mathmath333):

its not accepting \sigma , i have to hack it by some means

OpenStudy (mathmath333):

oh thats cool :D

OpenStudy (perl):

your solution is pretty cool, you used the identity sin^2 + cos^2 = 1 , but indirectly

OpenStudy (perl):

cos^2 = 1 - sin^2

OpenStudy (perl):

so either way you will end up at the same spot

OpenStudy (mathmath333):

yes but at first i thought it was wrong

OpenStudy (perl):

wolfram is a very forgiving math machine , it accepts a lot of forms

OpenStudy (perl):

i mean , it accepts a lot of different types of syntax

OpenStudy (mathmath333):

yes http://www.wolframalpha.com/input/?i=perl

OpenStudy (perl):

i see now, in your first post you subtracted the two equations , thats what those lines meant _ _ _ _ _ _

OpenStudy (perl):

:)

OpenStudy (perl):

wolfram is an encyclopedia?

OpenStudy (perl):

its a brief encyclopedia

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+\cdot \cdot \cdot \cdot \cdot+ \cos^2 90^{\circ} \hspace{.33em}\\~\\ =\sin^2 88^{\circ}+\sin^2 86^{\circ}+\sin^2 84^{\circ} \hspace{.33em}\\~\\ +\cdot \cdot \cdot \cdot \cdot+\sin^2 2^{\circ}+\sin^2 0^{\circ}-------\color{red}{(1)} \hspace{.33em}\\~\\ \cos^2 2^{\circ}+\cos^2 4^{\circ}+\cos^2 6^{\circ}+ \hspace{.33em}\\~\\ \cdot \cdot \cdot \cdot \cdot+\cos^2 90^{\circ} \hspace{.33em}\\~\\ =45-(\sin^2 2^{\circ}+\sin^2 4^{\circ}+\sin^2 6^{\circ}+ \hspace{.33em}\\~\\ \cdot \cdot \cdot \cdot \cdot+\sin^2 90^{\circ})-------\color{red}{(2)} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

those lines were these as there was not enough space in the page

OpenStudy (perl):

and then you added equation (1) and (2)

OpenStudy (mathmath333):

yes hah

OpenStudy (perl):

yeah thats nifty

OpenStudy (perl):

thats weird that wolfram was able to infer the pattern for cos 2 degrees + cos (4 degrees) + ... + cos ( 90 degrees) but for the square terms it was stumped

OpenStudy (perl):

want to do a problem that stumped wolfram, which has a simple solution?

OpenStudy (perl):

a trig equation

OpenStudy (perl):

solve sind(2x) = cosd(3x -10) sind and cosd only accepts arguments with degrees (so we dont have a problem with radians)

OpenStudy (mathmath333):

yes that happens wolfram is a machine and man made machine , machine did'nt made man so man>wolfram

OpenStudy (perl):

oh i guess it didn't stump it then,

OpenStudy (mathmath333):

once i had a problem where wolfram was wrong , but i don't remember

OpenStudy (perl):

hmm, i thought it did a week ago

OpenStudy (mathmath333):

lolol

OpenStudy (perl):

you can see the general solutions on the bottom http://www.wolframalpha.com/input/?i=sind+%282x%29%3Dcosd+%283x-10%29

OpenStudy (perl):

oh i remember what was the problem, i got a 'wild' solution from mathematica 10

OpenStudy (perl):

i will if i can attach the output

OpenStudy (perl):

attach file button is dead :/

OpenStudy (mathmath333):

haha http://www.wolframalpha.com/input/?i=who+are+you+%3F

OpenStudy (perl):

lol that didnt work

OpenStudy (mathmath333):

the file is not opening

OpenStudy (perl):

thats creepy

OpenStudy (perl):

the who are you

OpenStudy (perl):

its probably a joke

OpenStudy (perl):

LOL

OpenStudy (perl):

clever

OpenStudy (perl):

i uploaded the wild output using pdf, see if this work s

OpenStudy (perl):

i mean i saved as .pdf

OpenStudy (mathmath333):

what r u trying to do

OpenStudy (perl):

im trying to show you what mathematica 10 displays when i try to solve that equation above

OpenStudy (perl):

wolfram (web version) gives a more logical answer

OpenStudy (mathmath333):

oh k

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=tell+me+a+joke

OpenStudy (mathmath333):

every time it has a new joke lol,

OpenStudy (perl):

here http://www.filedropper.com/dd

OpenStudy (perl):

thats my file

OpenStudy (mathmath333):

that also isn;t working

OpenStudy (perl):

click on 'download this file'

OpenStudy (mathmath333):

ok but at the end it should interpret it as solve ------- for \(x\) it displaying \(\cos x ~and~\sin x\)

OpenStudy (perl):

were you able to open it ?

OpenStudy (mathmath333):

yes lol

OpenStudy (perl):

oh for x

OpenStudy (perl):

usually it solves for x by default

OpenStudy (perl):

if i say solve x^2 +2x + 1 = 0 , it will solve for x

OpenStudy (mathmath333):

u have to write it as if u wanr ,for integers , reals or complex

OpenStudy (perl):

i tried using 'over reals', didn't help. i dont know, mathematica 10 acts oddly. maybe i should go back to mathematica 9 , i liked that version

OpenStudy (perl):

i can try integer solutions

OpenStudy (perl):

ok now it seems to work when i stipulate integer solutions

OpenStudy (perl):

ok thanks a lot

OpenStudy (perl):

it looks like mathematica goes straight to the real 'exact' solutions, thats why there was such a long expression output

OpenStudy (kainui):

I would use this \[\Large \cos^2(2x)=\frac{1}{2}+\frac{1}{2}\cos(4x)\] and change the sum from \[\Large \sum_{n=1}^{45}\cos^2(2x)= \sum_{n=1}^{45}\frac{1}{2} + \frac{1}{2}\cos(4x)\] that way the cosine part goes from almost cos(0) to cos(180) so they will all cancel each other out.